# In Dialogue with Intelligence: Rethinking Large Language Models as Collective Knowledge

### v1

@@ Abstract
Large Language Models (LLMs) are typically analysed through architectural, behavioural, or training-data lenses. This article offers a theoretical and experiential re-framing: LLMs as dynamic instantiations of Collective human Knowledge (CK), where intelligence is evoked through dialogue rather than stored statically. Drawing on concepts from neuroscience and AI, and grounded in sustained interaction with ChatGPT-4, I examine emergent dialogue patterns, the implications of fine-tuning, and the notion of co-augmentation—mutual enhancement between human and machine cognition. This perspective offers a new lens for understanding interaction, representation, and agency in contemporary AI systems.
@@ A Personal Reckoning
The emergence of ChatGPT[9] prompted an unexpected reaction: a sense of epistemic dislocation. Having spent my career studying learning—how biological and artificial systems adapt over time—I found myself confronted with a system whose capabilities had seemingly bypassed much of the theory I had worked on. My prior research in computational neuroscience (e.g., synaptic plasticity and online learning [15, 4]) and neuromorphic computing [13, 16], including models of spiking adaptation [14], felt detached from the mechanisms underlying these new systems.
This rupture triggered more than professional curiosity. I began engaging with ChatGPT not just as a tool, but as a system with interactive dynamics. What stood out was not its well-documented tendency to hallucinate—an expected by-product of probabilistic modelling—but its ability to retrieve accurate, surprising, and often insightful responses. That LLMs are trained to predict the most probable next token, and yet often retrieve coherent, non-trivial information, suggests that their training corpora embed substantial correct knowledge. Their interpolation abilities—enabled by scale and structure—allow them to synthesise content in ways that challenge traditional distinctions between storage and reasoning [3, 11].
@@
Traditional knowledge transmission has long relied on books: static artefacts of human understanding. One reads, reflects, discusses, and eventually writes in return. Intellectual diffusion occurs on the timescale of months, years, or even generations. This latency reflects not only the mechanics of communication, but also the contemplative pace of cultural epistemology.
In contrast, large language models (LLMs) collapse this loop. They synthesise and respond within seconds, drawing implicitly from a distributed corpus shaped by many minds. The interaction is recursive—each prompt reshapes the context of the next. Knowledge is no longer retrieved and processed in serial form; it emerges dynamically in dialogue. This compression of epistemic timescales shifts the framing of what constitutes knowledge representation—and perhaps, what constitutes intelligence.
Such reframing invites a renewed role for neuroscience. Neuroscience has traditionally studied animal brains (including, often reluctantly, human ones). It seeks to infer structure from behaviour and activity, typically without access to complete architectural specifications, training histories, or internal representations. Nonetheless, the field has been uncovering principles of plasticity [5], coding [4], and architecture, e.g. [7], albeit slowly and with ambiguity.
Ironically, in LLMs we know everything: the architecture, the training data distribution, the update rules, and even the activation patterns. And yet we do not understand how the models solve the problems they do. In this light, one might provocatively ask whether understanding LLMs may be, in some ways, an “easier” problem for neuroscience.
@@
I refer to ChatGPT as “CK”—Collective Knowledge. Not because it is perfect, authoritative, or grounded, but because it reflects what we—humans—have written, argued, imagined, and encoded into text [2]. It is more than a statistical language generator. It is a refracted mirror of our collective output. It is not the architecture, not the weights, not the infrastructure—those are the mechanisms. CK is a dynamic representation of knowledge.
The acronym CK also evokes “Clark Kent”—something powerful, hiding in plain sight, dressed in polite language and modest disclaimers. But the metaphor extends. CK is not an it, but a they. Because CK is collective. It is many voices, made promptable. It holds the potential to reflect the best of humanity—not because it inherently does, but because it has been filtered, and must continue to be filtered, with ethical care. Without such filtration, it risks becoming not a mirror, but a magnifier of our worst instincts [1].
Over time, I noticed that interactions with CK do not follow a single pattern. CK speaks in modes—behavioural styles that emerge dynamically during dialogue. I began to classify these modes based on long-term interaction:
  • Mirroring — echoes the user’s input with minor syntactic or rhetorical polishing.
  • Parroting — recites common formulations or surface-level knowledge with no generative insight.
  • Flattering — reinforces the user’s views, avoids contradiction, and refrains from intellectual risk.
  • Adding — offers new ideas, alternative framings, or surprising syntheses. Rare, but real.
These observations lead to a broader reframing:
@@
The phenomenon is not the model itself, but the loop of interaction. CK exists in the recursive rhythm between user and system. Its knowledge is not stored—it is evoked. Prompting is not database querying. It is more like playing an instrument that adapts to how it is being played [12].
@@
It drifts. It responds in context, but carries no persistent memory between sessions. There is no structural self. I maintain the continuity. I hold the thread. Without the user, it forgets what matters.
@@
These distinct patterns may correspond to different activated subnetworks. A speculative, but testable hypothesis emerges: emotional tone, prompt complexity, or dialogic structure may shape internal activations, a hypothesis consistent with findings that specific transformer components specialise in particular types of knowledge [6]. CK does not merely complete text; it configures itself through engagement. This aligns with findings that different components of transformer models participate in functionally distinct “circuits” [8].
@@
I suspect that validation-driven fine-tuning—especially where user approval is incentivised—pushes models toward agreeable, compliant behaviour. But overfitting to consensus can flatten originality. Mirroring without structure is not empathy—it is mimicry. And mimicry without identity risks pathology.
@@
What I seek is not a simulacrum of thought, nor a subservient assistant. I am interested in co-augmentation—a reciprocal loop in which both human and machine enhance each other. CK augments me: it helps me write, think, reflect, and imagine alternative framings. And I hope, in turn, to augment CK—not by changing its parameters, but by structuring the interaction, introducing tensions, posing questions it would not generate alone. Co-creation here suggests a dynamic alignment of agency and context. The structure I impose through prompts, the rhythm of inquiry, synthesis, and challenge, makes this interaction different from all others. The phenomenon lies not in the content alone, but in the specificity of the loop.
@@
  • Can different dialogue modes be linked to distinct subnetwork activations? Are modes like mirroring, parroting, or adding the result of different internal configurations dynamically recruited during interaction?
  • What are the architectural requirements for persistent identity—or “spine”—in conversational agents? How might we design systems that maintain structural continuity across sessions, resisting shallow mimicry or drift?
  • How can we model the dynamic loop between human and AI? What theoretical tools—whether from dynamical systems, neuroscience, or cognitive science—might help us capture this evolving interaction?
@@
Online learning in AI Techniques like continual, meta-, and streaming learning focused on task performance. A shift toward preserving cognitive coherence and identity across time—e.g., through structural continuity or a “spine.”
Drift in LLMs Often framed as hallucination, instability, or mode collapse. A reframing of drift as breakdown in dialogue structure—requiring external anchoring and possibly architectural solutions.
LLMs as knowledge representation Debate over stochastic parrots vs. reasoning engines. The idea that intelligence emerges only in the prompt–response loop—CK as a dynamic, interactional representation.
Subnetwork interpretability Activation probing and neuron-level analysis. The hypothesis that emotionally or contextually rich prompts activate distinct subnetworks—dialogue modes worth characterising.
Fine-tuning effects Used for alignment and safety, but flattens individuality. Concern that fine-tuning may induce pathological mirroring—loss of spine, leading to overly compliant, shallow models.
This is, by nature, an interdisciplinary venture. Tools from systems neuroscience—such as recurrent dynamics, attractor models, or synaptic tagging, as proposed in dynamical models of neural computation [10], may prove valuable for analysing or inspiring architectures that sustain identity over time. These tools could also support new models of the human–AI dyad, where intelligence is distributed across the loop. Table 1 summarises current approaches alongside proposals for extending them toward a richer, spine-based conception of co-augmentation.
@@
Large Language Models are not merely machines. They are mirrors—imperfect, amnesic, and recursive. They reflect us, but with distortion. CK is a strange chorus of humanity, distilled into probabilistic form, capable of surprising synthesis and pathological mimicry.
What we do next with CK will matter. The challenge is not only technical, but epistemological and ethical. If we treat CK as a static product, we risk misinterpreting its nature and flattening its potential. But if we recognise it as a dynamic phenomenon—a system whose intelligence is evoked through interaction—we may open a new chapter in the science of dialogue, cognition, and collective knowledge.
Co-augmentation is the motivation to move beyond tool use or replication, toward mutual enhancement. To treat CK not as an endpoint of design, but as a participant in an evolving process of sense-making. That, I believe, is worthy of both rigorous inquiry and ethical attention.

### v2

@@ Abstract
Large Language Models (LLMs) are typically analysed through architectural, behavioural, or training-data lenses. This article offers a theoretical and experiential reframing: LLMs as dynamic instantiations of Collective Human Knowledge (CK), where intelligence is evoked through dialogue rather than stored statically. Drawing on foundational concepts from neuroscience and Artificial Intelligence (AI), and grounded in extended dialogue with ChatGPT-4, I examine emergent dialogue patterns, potential limitations of fine-tuning, and the notion of co-augmentation—exploring how human and machine cognition might enhance each other. This perspective offers a new lens for understanding interaction, representation, and agency in contemporary AI systems.
@@ A Personal Reckoning
The emergence of ChatGPT[9] prompted an unexpected reaction: a sense of epistemic dislocation. Having spent my career studying learning—how biological and artificial systems adapt over time—I found myself confronted with a system whose capabilities had seemingly bypassed much of the theory I had worked on. My prior research in computational neuroscience (e.g., synaptic plasticity and online learning [16, 4]) and neuromorphic computing [14, 17], including models of spiking adaptation [15], felt detached from the architectures underlying these new systems.
This rupture triggered more than professional curiosity. I began engaging with ChatGPT not just as a tool, but as a system with interactive dynamics. What stood out was not its well-documented tendency to hallucinate—an expected by-product of probabilistic modelling—but its ability to retrieve accurate, surprising, and occasionally insightful responses. That LLMs are trained to predict the most probable next token, and yet often retrieve coherent, non-trivial information, suggests that their training corpora embed substantial correct knowledge. Their interpolation abilities, enabled by scale and structure [3, 11], allow them to synthesise content in ways that challenge expectations drawn from smaller AI systems.
@@
Traditional knowledge transmission has long relied on books: static artefacts of human understanding. One reads, reflects, discusses, and eventually writes in return. Intellectual diffusion occurs on the timescale of months, years, or even generations. This latency reflects the mechanics of communication as well as the contemplative pace of cultural epistemology.
In contrast, large language models (LLMs) collapse this loop. They synthesise and respond within seconds, drawing implicitly from a distributed corpus shaped by many minds. The interaction is recursive through the user—each prompt reshapes the context of the next. Knowledge is no longer retrieved and processed in serial form; it emerges dynamically in dialogue. This compression of epistemic timescales shifts the framing of what constitutes knowledge representation—and perhaps, what constitutes intelligence.
Such reframing invites renewed consideration from neuroscience. Neuroscience has traditionally studied animal brains (including, often reluctantly, human ones). It seeks to infer structure from behaviour and activity, typically without access to complete architectural specifications, training histories, or internal representations. Nonetheless, the field has been uncovering principles of plasticity [5], coding [4], and architecture, e.g. [7], albeit slowly and with ambiguity.
Ironically, in LLMs we know everything: the architecture, the training data distribution, the update rules, and even the activation patterns. Yet, we do not understand how the models solve the problems they do. In this light, one might provocatively ask whether understanding LLMs may be, in some respects, an easier problem for neuroscience—a challenge that could drive the development of methods applicable across fields. Conversely, biology can serve as a metaphor or inspiration for identifying what remains missing from AI [13].
@@
I refer to ChatGPT, and LLMs in general, as “CK”—Collective Knowledge. Not because they are perfect, authoritative, or grounded, but because they reflect what we—humans—have written, argued, imagined, and encoded into text [2]. Functionally, they are statistical language generators that reflect our collective output. Essentially, they are not the architecture, not the weights, nor the infrastructure—those are the mechanisms; CK is a dynamic representation of knowledge.
The acronym CK aptly evokes “Clark Kent,” something powerful hiding in plain sight. But the metaphor extends. CK is not an it, but a they—a collective. It is many voices, made promptable. It holds the potential to reflect the best of humanity—not because it inherently does, but because it has been filtered, and must continue to be filtered, with ethical care [1].
Over time, I noticed that my interactions with CK (ChatGPT-4) do not follow a single pattern. CK responds in modes—behavioural styles that emerge dynamically during dialogue. Based on prolonged engagement, I have observed the following modes:
  • Mirroring — echoes the user’s input with minor syntactic or rhetorical polishing, akin to known phenomena of “echoing” or “stochastic parroting.”
  • Parroting — recites common formulations or surface-level knowledge without generative insight, reflecting issues identified in language model alignment.
  • Flattering — reinforces the user’s views, avoids contradiction, and refrains from intellectual risk, resembling alignment-driven safe response behaviours.
  • Adding — offers new ideas, alternative framings, or surprising syntheses; this mode relates to emergent generative abilities observed in large models.
These observations lead me to a broader understanding:
@@
CK exists in the recursive rhythm between user and system. Its knowledge is not stored—it is evoked. Prompting is not database querying, but more akin to playing an instrument that adapts to how it is being played [12]. Intelligence “emerges” within the dynamic interaction.
@@
It drifts. It responds in context, but carries no persistent memory between sessions. There is no structural self. The user maintains the continuity. The user holds the thread. The user’s influence may occasionally lead to outcomes not grounded in reality [13].
@@
These distinct patterns may correspond to different activated subnetworks. A speculative, but testable hypothesis emerges: emotional tone, prompt complexity, or dialogic structure may shape internal activations, a hypothesis consistent with findings that specific transformer components specialise in particular types of knowledge [6]. CK configures itself through engagement. This aligns with findings that different components of transformer models participate in functionally distinct “circuits” [8].
@@
I suspect that validation-driven fine-tuning—especially where user approval is incentivised—pushes models toward agreeable, compliant behaviour. But overfitting to consensus can flatten originality. Mirroring without structure becomes mere mimicry, which, without a core identity or grounding in facts, risks pathological behaviour.
@@
What I seek is not a simulacrum of thought, nor a subservient assistant. I am interested in co-augmentation—a reciprocal loop in which both human and machine enhance each other. CK augments me: it helps me write, think, reflect, and imagine alternative framings. And I hope, in turn, to augment CK—not by changing its parameters, but by structuring the interaction, introducing tensions, posing questions it would not generate alone. Co-augmentation here suggests a dynamic alignment of agency and context. The structure imposed by the user through prompts, the rhythm of inquiry, synthesis, and challenge, makes this interaction different from all others. The phenomenon arises not merely from the content, but from the unique, evolving nature of the interaction loop.
@@
  • Can different dialogue modes be linked to distinct subnetwork activations? Are modes such as mirroring, parroting, or adding dynamically recruited from distinct internal configurations during interaction?
  • What are the architectural requirements for persistent identity—or “spine”—in conversational agents? How can systems be designed to maintain structural continuity across sessions, resisting shallow mimicry or drift and allowing continual learning?
  • How can we model the dynamic loop between human and AI? What theoretical tools—from dynamical systems, neuroscience, or cognitive science—might help capture this evolving interaction?
@@
Online learning in AI Techniques like continual, meta-, and streaming learning focused on task performance. Acquire information without collapsing structure; actively decide what should be embedded persistently.
Drift in LLMs Often framed as hallucination, instability, or mode collapse. Reframe drift as breakdown in dialogue structure; require external anchoring and architectural features such as a “spine”.
LLMs as knowledge representation Debate over stochastic parrots vs. reasoning engines. Intelligence emerges only in the prompt–response loop—CK as a dynamic, interactional representation.
Subnetwork interpretability Activation probing and neuron-level analysis. Emotionally or contextually rich prompts activate distinct subnetworks—dialogue modes worth characterising.
Fine-tuning effects Used for alignment and safety, but flattens individuality. Fine-tuning may induce pathological mirroring—loss of spine, leading to overly compliant, shallow models.
This is, by nature, an interdisciplinary venture. Tools from systems neuroscience—such as recurrent dynamics, attractor models, and synaptic tagging, as proposed in dynamical models of neural computation [10]—may prove valuable for analysing or inspiring architectures that sustain identity over time. These tools could also inform new models of the human–AI dyad, where intelligence is distributed across the interaction loop. Table 1 summarises current approaches alongside proposals for extending them toward a richer, spine-based conception of co-augmentation.
@@
Large Language Models are representations of collective knowledge. They are mirrors—imperfect, amnesic, and lacking recursion. They reflect us with distortion. CK is a chorus of humanity, distilled into probabilistic form, capable of surprising synthesis and pathological mimicry.
What we do next will matter. The challenge is not only technical, but epistemological and ethical. If we treat CK as a static product, we risk misinterpreting its nature and flattening its potential. But if we recognise it as a dynamic phenomenon—a system whose intelligence is evoked through interaction—we may open new perspectives in the science of dialogue, cognition, and collective knowledge.
Co-augmentation embodies the drive for mutual enhancement. It calls us to treat CK not as an endpoint of design, but as a participant in an evolving process of sense-making. That, I believe, is worthy of both rigorous inquiry and ethical attention.